# Conditional PED-ANOVA: Hyperparameter Importance in Hierarchical & Dynamic Search Spaces

### v1

@@ Abstract.
We propose conditional PED-ANOVA (condPED-ANOVA), a principled framework for estimating hyperparameter importance (HPI) in conditional search spaces, where the presence or domain of a hyperparameter can depend on other hyperparameters. Although the original PED-ANOVA provides a fast and efficient way to estimate HPI within the top-performing regions of the search space, it assumes a fixed, unconditional search space and therefore cannot properly handle conditional hyperparameters. To address this, we introduce a conditional HPI for top-performing regions and derive a closed-form estimator that accurately reflects conditional activation and domain changes. Experiments show that naive adaptations of existing HPI estimators yield misleading or uninterpretable importance estimates in conditional settings, whereas condPED-ANOVA consistently provides meaningful importances that reflect the underlying conditional structure.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learning approachesccs: Mathematics of computing Mathematical analysis
@@ 1. Introduction
Analysis of parameter importance (Sobol, 1993, 2001; Hooker, 2007; Huang, 1998) is a central tool for understanding and designing optimization procedures. Over the past decade, techniques for measuring the importance of parameters have attracted considerable attention in the automated machine learning (AutoML) community (Hutter et al., 2014; Biedenkapp et al., 2017; Van Rijn and Hutter, 2018), particularly for improving the efficiency of hyperparameter optimization (HPO) (Bergstra et al., 2011; Bergstra and Bengio, 2012; Snoek et al., 2012; Lindauer et al., 2022). Given a hyperparameter search space and a set of evaluated configurations, methods for hyperparameter importance (or HPI methods for short) aim to quantify how strongly each hyperparameter influences the observed performance. Such information is useful for interpreting optimization runs, identifying critical design choices, detecting irrelevant or redundant parameters, and guiding subsequent search or manual tuning (Melis et al., 2018; Henderson et al., 2018). As a result, HPI is widely used across various settings, including HPO of machine learning models and HPO algorithm design, where understanding hyperparameter sensitivity is often essential (Van Rijn and Hutter, 2018; Zimmer et al., 2021; Probst et al., 2019; Theodorakopoulos et al., 2024).
@@
Classic HPI methods such as functional analysis of variance (f-ANOVA) (Hutter et al., 2014) quantify HPI by measuring how much variance in the objective can be attributed to each hyperparameter. However, their existing local variants mix contributions from unpromising regions, leading to inaccurate estimates of HPI in top-performing subspaces (Hutter et al., 2014; Biedenkapp et al., 2019). PED-ANOVA (Watanabe et al., 2023) mitigates this by defining local HPI within a top-performing subset of the search space and efficiently approximating it using the Pearson divergence between one-dimensional distributions of each hyperparameter. Moreover, PED-ANOVA computes both global and local HPIs substantially faster than f-ANOVA while maintaining accurate HPI estimates. Thanks to this combination of speed and reliability, it has been adopted for HPI analysis in practical HPO frameworks, such as Optuna (Akiba et al., 2019), demonstrating its utility in real-world workflows.
@@
However, we argue that PED-ANOVA has a practical limitation: it cannot properly handle hierarchical/dynamic search spaces, where the presence/domain of a hyperparameter depends on other hyperparameters (hereafter referred to as conditional search spaces). In practice, many real-world problems involve such conditional search spaces (Hutter et al., 2019). Typical examples include model selection, where hyperparameters are dependent on a chosen algorithm (Thornton et al., 2013; Kotthoff et al., 2017; Feurer et al., 2015; Komer et al., 2014), and neural network architecture design, where active parameters depend on the chosen layers (Bergstra et al., 2013; Swersky et al., 2014). Unfortunately, PED-ANOVA cannot be directly applied to such settings, since it assumes a fixed, unconditional search space and defines distributions over a single, fixed domain for each hyperparameter. We observe that naive extensions of PED-ANOVA, such as filtering inactive configurations, imputing missing values, or artificially expanding parameter domains, do not resolve this issue: these heuristics often violate the conditional semantics of the search space and yield unnatural or uninterpretable HPI estimates (cf. Section 5.4).
@@
To address this limitation, we propose conditional PED-ANOVA (condPED-ANOVA), a principled framework for estimating HPI in conditional search spaces. We identify a critical flaw in the original definition of local HPI in PED-ANOVA when applied to conditional search spaces: the importance of hyperparameters that induce conditionality is spuriously attributed to the hyperparameters whose domains or activation depend on them, leading to systematically misleading importance estimates. Consequently, merely constructing a method that correctly computes local HPI in conditional search spaces is still insufficient. We introduce conditional local HPI, which measures importance within top-performing regions without confounding from conditional activation or domain changes. Building on this definition, we derive a closed-form PED-ANOVA-style estimator that remains efficient and yields an interpretable HPI in conditional settings. Experiments confirm that condPED-ANOVA reliably produces a meaningful HPI that reflects the underlying conditional structure, while naive extensions of existing methods fail to do so.
@@
  • We introduce conditional local HPI, a definition that measures importance within top-performing regions while accounting for conditional activation and domain changes.
  • We derive a closed-form PED-ANOVA-style estimator for conditional local HPI, retaining the efficiency of the original PED-ANOVA via Pearson divergence of one-dimensional distributions estimated with KDE.
@@ 2.1. Variance-Based HPI and f-ANOVA
HPI aims to quantify how strongly each hyperparameter (and their interactions) influences performance, typically to support search-space design, debugging, and interpretation of HPO runs. Early work on parameter importance analysis builds on variance-based sensitivity analysis such as Sobol indices (Sobol, 1993, 2001) and ANOVA decompositions (Hooker, 2007; Huang, 1998), which attribute the global variance of an objective to individual coordinates and their interactions. In the context of HPO, f-ANOVA (Hutter et al., 2014) applies this idea to HPO objectives by fitting a surrogate model such as a random forest over the hyperparameter space and then estimating main and interaction effects by Monte Carlo integration over the surrogate. This yields importance values that reflect how much of the global variation in performance across the entire search space can be explained by each hyperparameter. f-ANOVA includes efficient marginal computation over random-forest predictions and produces a global ranking of important hyperparameters, making it a standard post-hoc analysis tool in AutoML and algorithm configuration.
@@ 2.2. PED-ANOVA
PED-ANOVA (Watanabe et al., 2023) extends HPI to settings where the region of interest is not an entire search space. It defines local HPI as the variance contribution of each hyperparameter within the top-$\gamma$ subset of the search space, thereby measuring importance directly in the most relevant region. It further shows that this local HPI can be approximated in closed form using the Pearson divergence between the one-dimensional distributions of each hyperparameter of two top-performing subsets, estimated via one-dimensional KDEs. Empirically, this leads to much faster HPI computation than f-ANOVA while maintaining accurate importance estimates in top-performing subspaces. The details of PED-ANOVA are reviewed in Section 3.
@@ 2.3. Conditional Hyperparameters in HPO
Many HPO algorithms furthermore operate on conditional search spaces, where certain hyperparameters are active only under specific parent choices, and their domains may also change depending on those choices (e.g., algorithm selection or internal mode switches). This view is made explicit in practical AutoML systems, such as Auto-WEKA (Thornton et al., 2013), auto-sklearn (Feurer et al., 2015, 2022), which describe conditional hyperparameters that are only relevant when a parent hyperparameter takes a particular value. Among widely used optimizers, SMAC (Lindauer et al., 2022) is designed to handle such conditional hierarchies in CASH-style spaces (Thornton et al., 2013), while TPE (Bergstra et al., 2011) formulates the search space as a tree-structured generative process where some variables are only well-defined under specific upstream choices. BOHB (Falkner et al., 2018) combines Hyperband (Li et al., 2018) with a Bayesian-optimization component built on TPE, thereby optimizing over the same tree-structured conditional spaces while exploiting cheap low-fidelity evaluations. Beyond these algorithms, a complementary line of work develops BO surrogates tailored to conditional domains to better share information across branches (Swersky et al., 2014; Jenatton et al., 2017; Ma and Blaschko, 2020).
@@ 2.4. HPI under Conditional Search Spaces
In conditional search spaces, applying standard HPI pipelines generally requires mapping configurations to a fixed representation, typically by imputing inactive hyperparameters with default values as done in SMAC-style toolchains and related preprocessing workflows (Lindauer et al., 2022). Practical analysis suites such as CAVE (Biedenkapp et al., 2019) (and its integration in BOAH (Lindauer et al., 2019)) then compute importance post-hoc using methods like f-ANOVA and local parameter importance, including for multi-fidelity runs. Optuna’s importance API highlights a common limitation of conditional spaces: by default, it evaluates only parameters that appear and have the same domain in all samples, excluding conditional parameters (Akiba et al., 2019). Recent work on multi-objective HPI likewise treats missing values arising from conditional hyperparameters in preprocessing before applying surrogate-based importance measures such as f-ANOVA and ablation paths (Theodorakopoulos et al., 2024).
@@ 3.2. Local HPI
$m_{\gamma}$ $\coloneqq\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\mid b_{\gamma}(\bm{x})]=\frac{\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\,b_{\gamma}(\bm{x})]}{\mathbb{E}_{\bm{x}\sim p_{0}}[b_{\gamma}(\bm{x})]},$
$f_{\gamma}^{(d)}(x^{(d)})$ $\coloneqq\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\mid b_{\gamma}(\bm{x}),x^{(d)}]$
$\,=\frac{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\,b_{\gamma}(\bm{x})\mid x^{(d)}]}{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[b_{\gamma}(\bm{x})\mid x^{(d)}]},$
@@
$v_{\gamma}^{(d)}$ $\coloneqq\mathrm{Var}_{\bm{x}\sim p_{0}}\bigl(f_{\gamma}^{(d)}(x^{(d)})\mid b_{\gamma}(\bm{x})\bigr)$
@@
$\text{Local HPI}\coloneqq v_{\gamma}^{(d)}\Big/\sum_{d^{\prime}=1}^{D}v_{\gamma}^{(d^{\prime})}.$
@@ 3.3. PED-ANOVA
$D_{\mathrm{PE}}(p\|q)\coloneqq\int\left(\frac{p(x)}{q(x)}-1\right)^{2}q(x)\mathrm{d}x.$
@@ 4.1. Regime-Based Representation of Conditional Hyperparameters
Algorithm 1 condPED-ANOVA Computation for the $d$-th Hyperparameter
@@
9:  $\alpha_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma^{\prime},\,i}|\,/\,|\mathcal{D}_{\gamma^{\prime}}|,\;\beta_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma,\,i}|\,/\,|\mathcal{D}_{\gamma}|$ $\triangleright$ regime probabilities in top $\gamma^{\prime}$ and $\gamma$ subsets
@@ 4.2. Conditional Local HPI via Within-Regime Variance
We observe that the standard local HPI based on the local marginal variance $v_{\gamma}^{(d)}$ mixes two distinct sources of variation: differences within each regime and differences between (inter) regimes. This can be seen through the following variance decomposition:
@@
$\hskip 25.00003pt+\underbrace{\mathrm{Var}_{I^{(d)}}\left(\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]\right)}_{\text{inter-regime variance}},$
@@
The inter-regime variance captures differences across regimes, weighted by their frequency in the top-$\gamma$ region. However, for conditional hyperparameters, the regime $I^{(d)}$ is determined by other hyperparameters (e.g., parent choices) rather than by the $d$-th hyperparameter itself. As a result, attributing this inter-regime effect to the $d$-th hyperparameter is often inappropriate: it can reflect changes in upstream decisions rather than the effect of the value of $x^{(d)}$ itself. In fact, we can show that the standard local marginal variance of a conditioned hyperparameter includes the conditioning variable’s local marginal variance, and is therefore always no smaller than it (cf. Theorem E.1 in Section E.1). In particular, this leakage implies that even inactive hyperparameters can be assigned non-negligible importance solely due to upstream choices. Such behavior is undesirable because tuning an inactive hyperparameter has no effect on the objective and is therefore meaningless.
Motivated by this consideration, we define the local HPI for conditional hyperparameters using only the within-regime component:
@@
We now derive a closed-form condPED-ANOVA estimator that efficiently computes the within-regime local marginal variance of each hyperparameter from one-dimensional PDFs within each regime. The conditional local HPI then follows immediately by normalization, as defined in Equation 3. The pseudocode is provided in Algorithm 1.
@@
$\alpha_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}|}{|\mathcal{D}_{\gamma^{\prime}}|},\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma,\,i}^{(d)}|}{|\mathcal{D}_{\gamma}|}.$
@@
With these definitions, we obtain the following closed-form expression for the within-regime local marginal variance:
@@
The proof of Theorem 4.2 is provided in Appendix B. Equation 5 aggregates within-regime distributional changes using regime-dependent weights $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$. Intuitively, this weighting emphasizes regimes that are enriched in the tighter top set (i.e., with larger $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$), while attributing importance only to changes within each regime. If a regime corresponds to an inactive configuration (so the value is always $\bot$), then $p_{\gamma^{\prime},\,i}^{(d)}$ and $p_{\gamma,\,i}^{(d)}$ are degenerate at $\bot$ and the corresponding divergence is zero. When the hyperparameter is not conditional, we have $K^{(d)}=1$ and $\alpha_{1}^{(d)}=\beta_{1}^{(d)}=1$, so Equation 5 reduces to the original PED-ANOVA in Equation 1.

### v2

@@ Abstract.
We propose conditional PED-ANOVA (condPED-ANOVA), a principled framework for estimating hyperparameter importance (HPI) in conditional search spaces, where the presence or domain of a hyperparameter can depend on other hyperparameters. Although the original PED-ANOVA provides a fast and efficient way to estimate HPI within the top-performing regions of the search space, it assumes a fixed, unconditional search space and therefore cannot properly handle conditional hyperparameters. To address this, we introduce a conditional HPI for top-performing regions and derive a closed-form estimator that accurately reflects conditional activation and domain changes. Experiments show that naive adaptations of existing HPI estimators yield misleading or uninterpretable importances in conditional settings, whereas condPED-ANOVA consistently provides meaningful importances that reflect the underlying conditional structure. Our code is publicly available at https://github.com/kAIto47802/condPED-ANOVA.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2026/02ccs: Computing methodologies Machine learning approachesccs: Mathematics of computing Mathematical analysis
@@ 1. Introduction
Analysis of parameter importance (Sobol, 1993, 2001; Hooker, 2007; Huang, 1998) is a central tool for understanding and designing optimization procedures. Over the past decade, techniques for measuring parameter importance have attracted considerable attention in the automated machine learning (AutoML) community (Hutter et al., 2014; Biedenkapp et al., 2017; van Rijn and Hutter, 2018), particularly for improving hyperparameter optimization (HPO) (Bergstra et al., 2011; Bergstra and Bengio, 2012; Snoek et al., 2012; Lindauer et al., 2022; Eggensperger et al., 2013). Given a hyperparameter search space and evaluated configurations, methods for hyperparameter importance (HPI) aim to quantify how strongly each hyperparameter influences the observed performance. Such information helps interpret optimization runs, identify critical design choices, detect irrelevant or redundant parameters, and guide subsequent search or manual tuning (Melis et al., 2018; Henderson et al., 2018). As a result, HPI is widely used across various settings, including HPO of machine learning models and HPO algorithm design, where understanding hyperparameter sensitivity is often essential (van Rijn and Hutter, 2018; Zimmer et al., 2021; Probst et al., 2019; Theodorakopoulos et al., 2024).
@@
Classic HPI methods such as functional analysis of variance (f-ANOVA) (Hutter et al., 2014) quantify HPI by measuring how much objective variance is attributed to each hyperparameter. However, their existing local variants mix contributions from unpromising regions, leading to inaccurate estimates of HPI in top-performing subspaces (Hutter et al., 2014; Biedenkapp et al., 2019). PED-ANOVA (Watanabe et al., 2023) addresses this by defining local HPI within a top-performing subset and efficiently approximating it using the Pearson divergence between one-dimensional distributions of each hyperparameter. Moreover, PED-ANOVA computes both global and local HPIs substantially faster than f-ANOVA while preserving accuracy. Thanks to this combination of speed and reliability, it has been adopted for HPI analysis in practical HPO frameworks, such as Optuna (Akiba et al., 2019), demonstrating its utility in real-world workflows.
@@
However, we argue that PED-ANOVA has a practical limitation: it cannot properly handle hierarchical/dynamic search spaces (hereafter conditional search spaces), where the presence or domain of a hyperparameter depends on others. In practice, many real-world problems involve conditional search spaces (Hutter et al., 2019). Typical examples include model selection, where hyperparameters depend on the chosen algorithm (Thornton et al., 2013; Kotthoff et al., 2017; Feurer et al., 2015; Komer et al., 2014), and neural network architecture design, where active parameters depend on the chosen layers (Bergstra et al., 2013; Swersky et al., 2014). Unfortunately, PED-ANOVA cannot be directly applied to such settings, as it assumes a fixed, unconditional search space and defines distributions over a single, fixed domain for each hyperparameter. We observe that naive extensions of PED-ANOVA, such as filtering inactive samples, imputing missing values, or artificially expanding parameter domains, do not resolve this issue: these heuristics often violate the conditional semantics of the search space and yield unnatural or uninterpretable HPI estimates (cf. Section 5.4).
@@
To address this limitation, we propose conditional PED-ANOVA (condPED-ANOVA), a principled framework for estimating HPI in conditional search spaces. We identify a critical flaw in PED-ANOVA’s original definition of local HPI under conditionality: the importance of hyperparameters that induce conditionality is spuriously attributed to dependent hyperparameters, yielding systematically misleading HPI estimates (empirically shown in Section 5.5.1 and theoretically in Theorem F.1). Consequently, merely constructing a method that correctly computes local HPI in conditional search spaces is still insufficient. We thus introduce conditional local HPI, which measures importance within top-performing regions without confounding from conditional activation or domain changes. Building on this definition, we derive a closed-form PED-ANOVA-style estimator that yields interpretable HPIs in conditional settings. Experiments confirm that condPED-ANOVA reliably produces a meaningful HPI that reflects the underlying conditional structure, while naive extensions of existing methods fail to do so.
@@
  • We introduce conditional local HPI, a definition that measures importance without being confounded by conditional activation and domain changes.
  • We derive a closed-form PED-ANOVA-style estimator, namely condPED-ANOVA, for conditional local HPI, retaining PED-ANOVA’s efficiency.
@@ 2.1. Variance-Based HPI and f-ANOVA
HPI aims to quantify how strongly each hyperparameter (and their interactions) influences performance, typically to support search-space design, debugging, and interpretation of HPO runs. Early work on parameter importance analysis builds on variance-based sensitivity analysis such as Sobol indices (Sobol, 1993, 2001) and ANOVA decompositions (Hooker, 2007; Huang, 1998), which attribute the global variance of an objective to individual coordinates and their interactions. In the context of HPO, f-ANOVA (Hutter et al., 2014) applies this idea to HPO objectives by fitting a surrogate model (e.g., a random forest) over the search space and estimating HPIs by integrating the surrogate over tree partitions using region-volume weights. This produces global importance scores that quantify how much of the performance variance is explained by each hyperparameter, and it has become a standard post-hoc analysis tool in AutoML.
@@ 2.2. PED-ANOVA
PED-ANOVA (Watanabe et al., 2023) extends HPI to settings where the region of interest is not an entire search space. It defines local HPI as the variance contribution of each hyperparameter within the top-$\gamma$ subset of the search space, thereby measuring importance directly in the most relevant region. It further shows that this local HPI can be approximated in closed form using the Pearson divergence between the one-dimensional distributions of each hyperparameter of the two top-performing subsets, estimated via one-dimensional kernel density estimators (KDEs) (Parzen, 1962). Empirically, this leads to much faster HPI computation than f-ANOVA while maintaining accurate importance estimates in top-performing subspaces. The details of PED-ANOVA are reviewed in Section 3.
@@ 2.3. Conditional Hyperparameters in HPO
Many HPO algorithms operate on conditional search spaces, where some hyperparameters’ activation and domains depend on specific parent choices (e.g., algorithm selection). This conditional structure is explicitly modeled in practical AutoML systems, such as Auto-WEKA (Thornton et al., 2013) and auto-sklearn (Feurer et al., 2015, 2022). Among widely used optimizers, SMAC (Lindauer et al., 2022) is designed to handle such conditional hierarchies in CASH-style spaces (Thornton et al., 2013), while TPE (Bergstra et al., 2011) formulates the search space as a tree-structured generative process. BOHB (Falkner et al., 2018) combines Hyperband (Li et al., 2018) with a TPE-based Bayesian optimization (BO), optimizing over the same tree-structured conditional spaces while exploiting cheap low-fidelity evaluations. Beyond these algorithms, a complementary work develops BO surrogates tailored to conditional domains to share information across branches (Swersky et al., 2014; Jenatton et al., 2017; Ma and Blaschko, 2020).
@@ 2.4. HPI under Conditional Search Spaces
In conditional search spaces, applying standard HPI pipelines requires mapping configurations to a fixed representation, often by imputing inactive hyperparameters with default values as done in SMAC-style toolchains (Lindauer et al., 2022). Practical analysis suites such as CAVE (Biedenkapp et al., 2019) (and its integration in BOAH (Lindauer et al., 2019)) and DeepCAVE (Segel et al., 2025) then compute importance post-hoc using methods like f-ANOVA and local parameter importance, including for multi-fidelity runs. Current Optuna’s importance API highlights a common limitation of conditional spaces: by default, it evaluates only parameters that appear and have the same domain in all samples, excluding conditional parameters (Akiba et al., 2019). Recent work on multi-objective HPI likewise treats missing values arising from conditional hyperparameters in preprocessing before applying surrogate-based importance measures such as f-ANOVA and ablation paths (Theodorakopoulos et al., 2024).
@@ 3.2. Local HPI
(1) $\hskip 27.11546ptm_{\gamma}\coloneqq\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\mid b_{\gamma}(\bm{x})]=\frac{\mathbb{E}_{\bm{x}\sim p_{0}}[f(\bm{x})\,b_{\gamma}(\bm{x})]}{\mathbb{E}_{\bm{x}\sim p_{0}}[b_{\gamma}(\bm{x})]},$
(2) $\begin{aligned} f_{\gamma}^{(d)}(x^{(d)})&\coloneqq\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\mid b_{\gamma}(\bm{x}),x^{(d)}]\\ &\,=\frac{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[f(\bm{x})\,b_{\gamma}(\bm{x})\mid x^{(d)}]}{\mathbb{E}_{\bm{x}^{(\bar{d}\,)}\sim p_{0}^{(\bar{d}\,)}}[b_{\gamma}(\bm{x})\mid x^{(d)}]},\end{aligned}$
@@
(3) $v_{\gamma}^{(d)}$ $\coloneqq\mathrm{Var}_{\bm{x}\sim p_{0}}\bigl(f_{\gamma}^{(d)}(x^{(d)})\mid b_{\gamma}(\bm{x})\bigr)$
@@
(4) $\text{Local HPI}\coloneqq v_{\gamma}^{(d)}\Big/\sum_{d^{\prime}=1}^{D}v_{\gamma}^{(d^{\prime})}.$
@@ 3.3. PED-ANOVA
(6) $D_{\mathrm{PE}}(p\|q)\coloneqq\int\left(\frac{p(x)}{q(x)}-1\right)^{2}q(x)\mathrm{d}x.$
@@ 4.1. Regime-Based Representation of Conditional Hyperparameters
Algorithm 1 condPED-ANOVA Computation for the $d$-th Hyperparameter $x^{(d)}$
@@
9:  $\alpha_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}|\,/\,|\mathcal{D}_{\gamma^{\prime}}|,\;\beta_{i}^{(d)}\leftarrow|\mathcal{D}_{\gamma,\,i}^{(d)}|\,/\,|\mathcal{D}_{\gamma}|$ $\triangleright$ regime probabilities in top $\gamma^{\prime}$ and $\gamma$ subsets
@@ 4.2. Conditional Local HPI via Within-Regime Variance
We observe that the standard local HPI based on the local marginal variance $v_{\gamma}^{(d)}$ mixes two distinct sources of variation: differences within each regime and differences between (inter) regimes. This can be seen through the following variance decomposition by the law of total variance:
@@
$\hskip 26.00009pt+\underbrace{\mathrm{Var}_{I^{(d)}}\left(\mathbb{E}_{Z^{(d)}}\left[g_{\gamma}^{(d)}(I^{(d)},Z^{(d)})\,\middle|\,I^{(d)}\right]\right)}_{\text{inter-regime variance}},$
@@
The inter-regime variance captures differences across regimes, weighted by their frequency in the top-$\gamma$ region. However, for conditional hyperparameters, the regime $I^{(d)}$ is determined by other hyperparameters (e.g., parent choices) rather than by the $d$-th hyperparameter itself. Thus, attributing this inter-regime effect to the $d$-th hyperparameter is often inappropriate. In fact, we can show that the standard local marginal variance of a conditioned hyperparameter is lower-bounded by the conditioning variable’s local marginal variance, implying systematic leakage (cf. Theorem F.1 in Section F.1). Consequently, even inactive hyperparameters can receive non-negligible importance solely due to upstream choices. Such behavior is undesirable because tuning an inactive hyperparameter has no effect on the objective and is therefore meaningless.
Motivated by this consideration, we define the local HPI for conditional hyperparameters using only the within-regime component.
@@
We now derive a closed-form condPED-ANOVA estimator that efficiently computes the within-regime local marginal variance of each hyperparameter. We adopt a PED-ANOVA-style construction since it enables fast HPI computation while maintaining accurate estimates (cf. runtime comparisons in Section D.1). The conditional local HPI then follows immediately by normalization, as defined in Equation 8. The pseudocode is provided in Algorithm 1.
@@
(10) $\alpha_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma^{\prime},\,i}^{(d)}|}{|\mathcal{D}_{\gamma^{\prime}}|},\quad\text{and}\quad\beta_{i}^{(d)}\coloneqq\frac{|\mathcal{D}_{\gamma,\,i}^{(d)}|}{|\mathcal{D}_{\gamma}|}.$
@@
With these definitions, we obtain the following closed-form estimation for the within-regime local marginal variance.
@@
The proof is provided in Appendix B. Equation 11 aggregates within-regime Pearson divergence using regime-dependent weights $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$. Intuitively, this weighting emphasizes regimes that are enriched in the tighter top set (i.e., with larger $\big(\alpha_{i}^{(d)}\big)^{2}/\beta_{i}^{(d)}$), while attributing importance only to changes within each regime. If a regime corresponds to an inactive configuration (so the value is always $\bot$), then $p_{\gamma^{\prime},\,i}^{(d)}$ and $p_{\gamma,\,i}^{(d)}$ are degenerate at $\bot$ and the corresponding divergence is zero. When the hyperparameter is not conditional, we have $K^{(d)}=1$ and $\alpha_{1}^{(d)}=\beta_{1}^{(d)}=1$, so Equation 11 reduces to the original PED-ANOVA in Equation 5.
5. Experiments on Synthetic Problems